OpenStudy (anonymous):

determine value h(2a) for function h(x)=x-1/1-x

OpenStudy (anonymous):

i got 2a-1/1-2a=H(x) do i need to do anything else

OpenStudy (anonymous):

its (2a-1)/(1-2a) =-1 , so -1 is the answer

OpenStudy (unklerhaukus):

here are the steps akash_809 has used \[h(x)=\frac{x-1}{1-x}\]\[h(2a)=\frac{2a-1}{1-2a}\]\[\qquad\quad=\frac{-(1-2a)}{1-2a}\]\[\qquad\quad=\frac{-\cancel{(1-2a)}}{\cancel{1-2a}}\]\[\qquad\quad =-1\]

OpenStudy (anonymous):

what if I used A+1 as my x value?

OpenStudy (anonymous):

platinum ring again :)

OpenStudy (unklerhaukus):

ring ring; hello?

OpenStudy (anonymous):

@jethro210 it would be same

OpenStudy (anonymous):

i.e =-1

OpenStudy (unklerhaukus):

the function gives the same value for all inputs (except x=1, which has indeterminate value )

OpenStudy (anonymous):

thanks for the medal un(c)kle

OpenStudy (unklerhaukus):

anytime

OpenStudy (anonymous):

@jethro210 u wl always get quick replies due to the pic :) you have used

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

My wife and I. THis site is awesome. I have been out of school for long time just decided to go back, found lots of help here